# Black-Hole Cartography

### v1

@@ II.3 Mapping
The spatial mapping technique turns the mode mixing coefficients from a fixed set of numbers imposed on the QNM model to a set of parameters that can be freely varied by the least-squares fitting algorithm.
To implement this, the QNM model in Eq. II.1 can be modified as follows,
@@
Here, $\tilde{\alpha}$ is mode that is being spatially mapped. The first summation is the same as in the usual model in Eq. 3 and includes all modes except the one that is mapped. The second summation has only one QNM, the mapped mode $\tilde{\alpha}$, but includes an independent copy of it separately in each spherical harmonic. In this model, $C_{\alpha}$ and $A^{\ell m}_{\tilde{\alpha}}$ are regarded as free complex amplitudes to be determined from the NR data by least-squares fitting. The mapping model in Eq. 11 can be expanded in spherical harmonics using the mode mixing coefficients in Eq. 4,
@@
In this form it is more easily seen how the new free amplitude parameters $A^{\ell m}_{\tilde{\alpha}}$ take on the role of the mixing coefficients for the mapped mode $\tilde{\alpha}$.
@@
It is also possible to extend this approach to map more than one QNM at the same time by removing another term, say $\tilde{\alpha}^{\prime}$, from the first sum in Eq. 11 and including another sum over $\ell,m$ for the new mapped mode $\tilde{\alpha}^{\prime}$ with coefficients $A^{\ell m}_{\tilde{\alpha}^{\prime}}$. However, the quality of the reconstruction (e.g. as quantified using $\mathcal{M}_{\tilde{\alpha}}$) decreases as the number of mapped modes is increased. An example of this is shown in Appendix B. The details of the changes to the least squares fitting procedure needed for the spatial mapping are detailed in Appendix A.
@@ II.4 CCE waveforms
Figure 2: Mismatches for linear QNM overtone model fits to the NR simulation 0001. In both panels, the gray region indicates the NR error estimate described in Appendix C. When a curve first dips below the numerical error, the line style changes to dashed. Left: The individual mode mismatch in the $\ell=m=2$ spherical harmonic mode. The fit uses a QNM model with fundamental plus overtones, including the $\alpha=(2,2,n,+)$ modes with $n\in\{0,1,\ldots,n_{\rm max}\}$. Right: The angle-averaged mismatch calculated across all (thick lines) $\ell,m$ spherical harmonic modes up to $\ell_{\rm max}=8$. The model includes all QNMs up to those with $\ell_{\rm max}=8$ and $n=n_{\rm max}$. As with the single mode case, increasing the number of overtones reduces the overall mismatch and shifts the minima of the mismatch to earlier times, associated with an earlier starting time for the ringdown, but it does so more smoothly in the angle-averaged case. A significantly improved angle-averaged match is achieved if the axisymmetric $m=0$ spherical harmonics are excluded from this fit (thin lines). These modes are poorly fit by QNMs.
@@
Figure 3: The spatial mismatch between spatially reconstructed QNMs and the expected spheroidal harmonics. Left: The spatial mismatch for the loudest QNM for each $\ell$ in simulation 0001 up to $\ell=6$. Because $q=1$, the loudest mode in the $\ell=3$ sector is the $m=2$ mode; similarly for the other odd $\ell$ values. The grey region indicates the NR error in the (2,2,0,+) mode mapping, described in Appendix C. The other modes have different (larger) associated errors which are not shown here, but the change in the style of each from solid to dashed indicates the point where the spatial mismatches first drop below their corresponding error thresholds. Right: The spatial mismatch for the overtones $\tilde{\alpha}=(2,2,n_{\rm max},+)$ in simulation 0001. The gray region indicates the NR error in the $n=0$ QNM, and the solid to dashed transition in the other lines indicates the point where the spatial mismatches first drops below its corresponding error threshold. The first three overtones are shown and their reconstructions get progressively worse but show minima in the spatial mismatches at earlier start times, consistent with their faster decay rates. The higher order overtones have worse spatial mismatches and are not shown.
@@
This work uses NR waveforms from the Spectral Einstein Code (SpEC), found in the Simulating Extreme Spacetimes (SXS) catalog [50, 49]. Analysis was performed with the 13 publicly available simulations with waveforms obtained using Cauchy-characteristic extraction (CCE). These waveform data were created using the SpECTRE CCE module [51, 52, 53].
@@
Currently, only a limited number of CCE waveforms are available in the public catalog; see Table 1. These simulations sparsely cover a range of progenitor parameters such as spin magnitudes, orientations, and binary mass ratio, $q$. All simulations were transformed into the superrest frame at $300M$ after the peak strain using the scri package. The waveforms were subsequently shifted in time so that the peak of the $22$ mode amplitude, $|\mathfrak{h}^{22}(t)|$, occurs at $t=0$ [54, 55, 56, 40].
@@
For each simulation taken from the catalog, the highest resolution level and the second smallest worldtube radius are together used as the preferred CCE waveform. Waveforms from other levels and worldtube radii were used to estimate the size of the NR errors, as described in Appendix C.
@@ III Linear QNM results
The new spatial mapping methods are demonstrated by first applying them to linear QNMs. In this case, first-order PT makes a clear, unambiguous prediction for the shape of the QNMs – the spin-weight $s=-2$ spheroidal harmonics. This provides a way to test the BH cartography programme.
@@
Figure  2 shows the mismatch results for simple QNM fits to simulation 0001. The left panel shows the mismatch in the $\ell=m=2$ spherical mode when fitted using a QNM model with a variable number of overtones. This is similar to Fig. 1 in Ref [27]. The right panel shows the angle-average mismatch calculated using all modes when fitted using a QNM model including all the QNMs (up to $\ell_{\rm max}$ and $n_{\rm max}$).
@@
Both panels in Fig. 2 display a similar structure, starting with a rapidly declining mismatch approaching a minimum at a time on or after $t_{0}=0$. At later times, the mismatch increases as the signal decays. The minima get pushed to earlier start times with an increasing number of overtones, related to the possibility of using an earlier choice of start time. The start times that give the minimum angle-average mismatch for the $n_{\mathrm{max}}=7$ QNM model are listed for each simulation in Table 1; these start times are used in subsequent figures.
@@
The accuracy of the angle-averaged QNM fits are limited by the $m=0$ spherical harmonic modes. In the right panel of Fig. 2, the thin lines indicate the angle-average mismatch with $m=0$ modes excluded from the fit. There is a considerable improvement in the mismatches, down to the level expected given the numerical noise present in the simulation. This shows that the $m=0$ are fit less well by the QNM model than all the other $m$ modes. This may be due to the presence of the non-oscillatory GW memory signal in these modes. There may also be an imperfect mapping to the superrest frame for these modes, affecting fits at late times.
@@
Having established that the QNM models are able to faithfully represent the NR data, we now turn to mapping. Figure 3 shows the results of spatially mapping various QNMs. The left panel shows the spatial mismatch of some of the loudest fundamental ($n=0$) QNMs in simulation 0001. Because this is a simulation of equal-mass quasicircular binary BHs ($q=1$), the symmetry $\phi\rightarrow\phi+\pi$ forces all spherical harmonic modes with odd $m$ indices to be identically zero. Therefore, the loudest mode in, for example, the $\ell=3$ sector is the $(\ell,m)=(3,2)$ mode. The right panel shows the spatial mismatch for some of the $\ell=m=2$ QNM overtones, with the largest overtone in the model being mapped in each case, analogously to the right panel in Fig. 2. In both panels, the mismatch of the fundamental mode is considerably smaller than that of the subdominant modes. Generally, the mismatch values decrease with increasing amplitude of the QNM.
@@
In the figures above, the mismatches are plotted as a function of start time. If the mode reconstruction is performed with a different choice of start time, then the values obtained for the $A^{\ell m}_{\tilde{\alpha}}$ amplitudes are different. Typically, for a reasonable range of $t_{0}$ values in the ringdown, the magnitude of the amplitude decreases exponentially and the phases of the amplitudes increase linearly with increasing $t_{0}$. The decreasing amplitude is due to the exponential decay of the QNM with a timescale given by the reciprocal of imaginary part of $\omega_{\tilde{\alpha}}$. The changing phase corresponds to the rotation of the perturbation sourcing the QNM around the symmetry axis of the BH. The rotation occurs at a frequency related to the real part of $\omega_{\tilde{\alpha}}$. To track this rotation, the following angle can be used,
@@
Figure 4: The top four figures show the real part of the spatial reconstruction of the (2,2,0,+) QNM at the start times indicated by squares on the main plot. These demonstrate the rotation of the QNM around the sphere. Main plot: Derivative of the argument $\Phi_{\alpha}$ of the spatial mismatch (corresponding to the phase shift between the mapped and predicted spatial structure) with respect to ringdown start time $t_{0}$. The solid lines show this quantity for simulation 0001 for the same 5 QNMs (matching colors) used in the left panel of Fig. 3. The shaded regions give an estimate of the error in the reconstructed frequencies and are described in Appendix C. The horizontal dashed lines give the predicted rotation frequencies, which are the real part of the complex QNM frequencies, $\omega_{\tilde{\alpha}}$. An animated version like these showing the rotation around the BH is available at [57].
@@
Figure 5: The magnitude of the decay-corrected amplitudes, $|\hat{A}^{\ell m}_{(220+)}|$, obtained from spatial mapping of the (2,2,0,+) mode for the 0001 waveform are plotted. The thick colored lines show the spherical modes indicated in the legend that are expected to have large amplitude. Translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 2$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the ratios of the spheroidal harmonics predicted by linear PT (expressed in terms of ratios of mode mixing coefficients) are plotted as horizontal dashed lines. The translucent area surrounding the lines on the three panels indicates the NR uncertainty in the values obtained. The way this is determined is described in Appendix C.
The $m\neq 2$ modes have also been included as translucent lines. In a multimode fit, all mixing coefficients where $m\neq 2$ would be zero and therefore so would the corresponding amplitudes. In contrast, here, where all spherical modes have been included in the spatial mapping, these amplitudes appear as non-zero values. The translucent gray lines moving diagonally along the figure are vanishing spherical modes, which are being scaled by the decay correction absorbed into the amplitude. The remaining translucent red lines are spherical modes for which the spatial mapping has attributed some non-zero amplitude. The contrast in the structure of these modes and the $m=2$ modes suggests that spatial mapping is correctly identifying the relevant spatial structure of the $(2,2,0,+)$ mode. The non-zero amplitudes of the $m\neq 2$ modes can be considered an indication of the numerical accuracy of the spatial mapping technique.
The bottom panels in Fig. 5 shows the magnitude of the ratios of the amplitudes in the $(2,2)$ and either $(3,2)$ or $(4,2)$ spherical modes respectively. These values agree with the magnitudes of the ratios of the corresponding mixing coefficients, which are also plotted. This provides an alternative mode-specific way to verify the accuracy of the spatial mapping technique beyond using the spatial mismatch. This is particularly helpful in checking that the reconstruction is not simply a consequence of overfitting to the loudest spherical mode (which, in this case, would be $\ell=m=2$).
Figure 6: The spatial mismatch for the best three reconstructed modes for all CCE waveforms. Highlighted columns indicate precessing simulations. In all cases, it was possible to reconstruct at least three QNMs with a spatial mismatch of $\lesssim 10^{-5}$ with their associated spheroidal harmonics. In all cases, except the precessing and asymmetric 0013 simulation, the best mapped mode is the fundamental $(2,2,0,+)$ QNM. The identity of the next best mapped QNMs changes depending on the parameters of the progenitor binary. In all cases, the best three mapped QNMs are all prograde $n=0$ modes.
@@
The analysis in this section demonstrates that spatial mapping can effectively reconstruct the spatial structure of QNMs without explicitly providing mode mixing coefficients. At first-order, where PT makes clear predictions about the spatial structure, the technique offers another method for detecting QNMs and may, in analogy with the single and multi-mode fits in Fig. 2, provide another way to determine the start time of the ringdown.
Figure 7: Leftmost panel: The real part of the normalized spatial reconstruction of the dominant QQNM with $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ for the 0001 NR simulation. Other panels: The various predictions for the shape of this QQNM discussed in the main text. Prediction A clearly has qualitatively the wrong shape and is therefore discarded. The remaining B, C, and D predictions all have qualitatively similar shapes (from B(i) to D, the spatial mismatches with the reconstruction are $\mathcal{M}_{\tilde{\alpha}}=7.4\times 10^{-4}$, $\mathcal{M}_{\tilde{\alpha}}=2.9\times 10^{-4}$, $\mathcal{M}_{\tilde{\alpha}}=2.5\times 10^{-4}$, and $\mathcal{M}_{\tilde{\alpha}}=2.6\times 10^{-4}$) and are further investigated in Fig. 8.

### v2

@@ II.3 Mapping
The spatial mapping technique turns the mode-mixing coefficients from a fixed set of numbers imposed on the QNM model to a set of parameters that can be freely varied by the least-squares fitting algorithm.
To implement this, the QNM model in Eq. 2 can be modified as follows,
@@
Here, $\tilde{\alpha}$ is the mode that is being spatially mapped. The first summation is the same as in the usual model in Eq. 3 and includes all modes except the one that is mapped. The second summation has only one QNM, the mapped mode $\tilde{\alpha}$, but includes an independent copy of it separately in each spherical harmonic. In this model, $C_{\alpha}$ and $A^{\ell m}_{\tilde{\alpha}}$ are regarded as free complex amplitudes to be determined from the NR data by least-squares fitting. The mapping model in Eq. 11 can be expanded in spherical harmonics using the mode-mixing coefficients in Eq. 4,
@@
In this form, it is more easily seen how the new free amplitude parameters $A^{\ell m}_{\tilde{\alpha}}$ take on the role of the mixing coefficients for the mapped mode $\tilde{\alpha}$.
@@
It is also possible to extend this approach to map more than one QNM at the same time by removing another term, say $\tilde{\alpha}^{\prime}$, from the first sum in Eq. 11 and including another sum over $\ell,m$ for the new mapped mode $\tilde{\alpha}^{\prime}$ with coefficients $A^{\ell m}_{\tilde{\alpha}^{\prime}}$. However, the quality of the reconstruction (e.g. as quantified using $\mathcal{M}_{\tilde{\alpha}}$) decreases as the number of mapped modes is increased. An example of this is shown in Appendix C. The details of the changes to the least-squares fitting procedure needed for the spatial mapping are detailed in Appendix A.
@@ II.4 CCE waveforms
Figure 2: Mismatches for linear QNM overtone model fits to the NR simulation 0001. In both panels, the gray region indicates the NR error estimate described in Appendix D. When a curve first dips below the numerical error, the line style changes to dashed. Left: The individual mode mismatch in the $\ell=m=2$ spherical harmonic mode. The fit uses a QNM model with fundamental plus overtones, including the $\alpha=(2,2,n,+)$ modes with $n\in\{0,1,\ldots,n_{\rm max}\}$. Right: The angle-averaged mismatch calculated across all (thick lines) $\ell,m$ spherical harmonic modes up to $\ell_{\rm max}=8$. The model includes all QNMs up to those with $\ell_{\rm max}=8$ and $n=n_{\rm max}$. As with the single-mode case, increasing the number of overtones reduces the overall mismatch and shifts the minima of the mismatch to earlier times, associated with an earlier starting time for the ringdown, but it does so more smoothly in the angle-averaged case. A significantly improved angle-averaged match is achieved if the axisymmetric $m=0$ spherical harmonics are excluded from this fit (thin lines). These modes are poorly fit by QNMs.
@@
Figure 3: The spatial mismatch between spatially reconstructed QNMs and the expected spheroidal harmonics. Left: The spatial mismatch for the loudest QNM for each $\ell$ in simulation 0001 up to $\ell=6$. Because $q=1$, the loudest mode in the $\ell=3$ sector is the $m=2$ mode similar for the other odd $\ell$ values. The grey region indicates the NR error in the (2,2,0,+) mode mapping, described in Appendix D. The other modes have different (larger) associated errors which are not shown here, but the change in the style of each from solid to dashed indicates the point where the spatial mismatches first drop below their corresponding error thresholds. Right: The spatial mismatch for the overtones $\tilde{\alpha}=(2,2,n_{\rm max},+)$ in simulation 0001. The gray region indicates the NR error in the $n=0$ QNM, and the solid to dashed transition in the other lines indicates the point where the spatial mismatch first drops below its corresponding error threshold. The first three overtones are shown and their reconstructions get progressively worse but show minima in the spatial mismatches at earlier start times, consistent with their faster decay rates. The higher-order overtones have worse spatial mismatches and are not shown.
@@
This work uses NR waveforms from the Spectral Einstein Code (SpEC), found in the Simulating Extreme Spacetimes catalog [51, 50]. Analysis was performed with the 13 publicly available simulations with waveforms obtained using CCE. These waveform data were created using the SpECTRE CCE module [52, 53, 54].
@@
Currently, only a limited number of CCE waveforms are available in the public catalog (see Table 1). These simulations sparsely cover a range of progenitor parameters, such as spin magnitudes, orientations, and the binary mass ratio, $q$. All simulations were transformed into the superrest frame at $300M$ after the peak strain using the scri package. The waveforms were subsequently shifted in time so that the peak of the $22$ mode amplitude, $|\mathfrak{h}^{22}(t)|$, occurs at $t=0$ [55, 56, 57, 40].
@@
For each simulation taken from the catalog, the highest resolution level and the second smallest worldtube radius were together used as the preferred CCE waveform. Waveforms from other levels and worldtube radii were used to estimate the size of the NR errors, as described in Appendix D.
@@ III Linear QNM results
The new spatial mapping methods are demonstrated by first applying them to linear QNMs. In this case, first-order PT makes a clear, unambiguous prediction for the shape of the QNMs – the spin-weight $s=-2$ spheroidal harmonics. This provides a way to test the BH cartography program.
@@
Figure  2 shows the mismatch results for simple QNM fits to simulation 0001. The left panel shows the mismatch in the $\ell=m=2$ spherical mode when fitted using a QNM model with a variable number of overtones. This is similar to Fig. 1 in Ref [27]. The right panel shows the angle-averaged mismatch calculated using all modes when fitted using a QNM model including all the QNMs (up to $\ell_{\rm max}$ and $n_{\rm max}$).
@@
Both panels in Fig. 2 display a similar structure, starting with a rapidly declining mismatch approaching a minimum at a time on or after $t_{0}=0$. At later times, the mismatch increases as the signal decays. The minima get pushed to earlier start times with an increasing number of overtones, related to the possibility of using an earlier choice of start time. The start times that give the minimum angle-averaged mismatch for the $n_{\mathrm{max}}=7$ QNM model are listed for each simulation in Table 1; these start times are used in subsequent figures.
@@
The accuracy of the angle-averaged QNM fits are limited by the $m=0$ spherical harmonic modes. In the right panel of Fig. 2, the thin lines indicate the angle-averaged mismatch with $m=0$ modes excluded from the fit. There is a considerable improvement in the mismatches, down to the level expected given the numerical noise present in the simulation. This shows that the $m=0$ are fit less well by the QNM model than all the other $m$ modes. This may be due to the presence of the nonoscillatory GW memory signal in these modes. There may also be an imperfect mapping to the superrest frame for these modes, affecting fits at late times.
@@
Having established that the QNM models are able to faithfully represent the NR data, we now turn to mapping. Figure 3 shows the results of spatially mapping various QNMs. The left panel shows the spatial mismatch of some of the loudest fundamental ($n=0$) QNMs in simulation 0001. Because this is a simulation of equal-mass quasicircular binary BHs ($q=1$), the symmetry $\phi\rightarrow\phi+\pi$ forces all spherical harmonic modes with odd $m$ indices to be identically zero. Therefore, the loudest mode in, for example, the $\ell=3$ sector is the $(\ell,m)=(3,2)$ mode. The right panel shows the spatial mismatch for some of the $\ell=m=2$ QNM overtones, with the largest overtone in the model being mapped in each case, analogously to the right panel in Fig. 2. In both panels, the mismatch of the fundamental mode is considerably smaller than that of the subdominant modes. Generally, the mismatch values decrease with the increasing amplitude of the QNM.
@@
In the figures above, the mismatches are plotted as a function of the start time. If the mode reconstruction is performed with a different choice of start time, then the values obtained for the $A^{\ell m}_{\tilde{\alpha}}$ amplitudes are different. Typically, for a reasonable range of $t_{0}$ values in the ringdown, the magnitude of the amplitude decreases exponentially, and the phases of the amplitudes increase linearly with increasing $t_{0}$. The decreasing amplitude is due to the exponential decay of the QNM with a timescale given by the reciprocal of the imaginary part of $\omega_{\tilde{\alpha}}$. The changing phase corresponds to the rotation of the perturbation sourcing the QNM around the symmetry axis of the BH. The rotation occurs at a frequency related to the real part of $\omega_{\tilde{\alpha}}$. To track this rotation, the following angle can be used,
@@
Figure 4: The top four figures show the real part of the spatial reconstruction of the (2,2,0,+) QNM at the start times indicated by squares on the main plot. These demonstrate the rotation of the QNM around the sphere. Main plot: Derivative of the argument $\Phi_{\alpha}$ of the spatial mismatch (corresponding to the phase shift between the mapped and predicted spatial structures) with respect to ringdown start time $t_{0}$. The solid lines show this quantity for simulation 0001 for the same 5 QNMs (matching colors) used in the left panel of Fig. 3. The shaded regions give an estimate of the error in the reconstructed frequencies and are described in Appendix D. The horizontal dashed lines give the predicted rotation frequencies, which are the real part of the complex QNM frequencies, $\omega_{\tilde{\alpha}}$. An animated version of these outcomes showing the rotation around the BH is available at [58].
@@
Figure 5: The magnitudes of the decay-corrected amplitudes, $|\hat{A}^{\ell m}_{(220+)}|$, obtained from spatial mapping of the (2,2,0,+) mode for the 0001 waveform are plotted. The thick colored lines show the spherical modes indicated in the legend that are expected to have large amplitudes. Translucent grey lines indicate odd $m$ modes, and the translucent red lines indicate even $m\neq 2$ modes. The ratios of selected amplitudes are plotted in the lower panels. The values for the ratios of the spheroidal harmonics predicted by linear PT (expressed in terms of ratios of mode-mixing coefficients) are plotted as horizontal dashed lines. The translucent area surrounding the lines on the three panels indicates the NR uncertainty in the values obtained. The way this is determined is described in Appendix D.
The $m\neq 2$ modes have also been included as translucent lines. In a multimode fit, all mixing coefficients where $m\neq 2$ would be zero and, therefore, so would the corresponding amplitudes. In contrast, here, where all spherical modes have been included in the spatial mapping, these amplitudes appear as nonzero values. The translucent gray lines moving diagonally along the figure are vanishing spherical modes, which are being scaled by the decay correction absorbed into the amplitude. The remaining translucent red lines are spherical modes for which the spatial mapping has attributed some nonzero amplitude. The contrast in the structure of these modes and the $m=2$ modes suggests that spatial mapping is correctly identifying the relevant spatial structure of the $(2,2,0,+)$ mode. The nonzero amplitudes of the $m\neq 2$ modes can be considered an indication of the numerical accuracy of the spatial mapping technique.
The bottom panels in Fig. 5 show the magnitude of the ratios of the amplitudes in the $(2,2)$ and either $(3,2)$ or $(4,2)$ spherical modes respectively. These values agree with the magnitudes of the ratios of the corresponding mixing coefficients, which are also plotted. This provides an alternative mode-specific way to verify the accuracy of the spatial mapping technique beyond using the spatial mismatch. This is particularly helpful in checking that the reconstruction is not simply a consequence of overfitting to the loudest spherical mode (which, in this case, would be $\ell=m=2$).
Figure 6: The spatial mismatch for the best three reconstructed modes for all CCE waveforms. Highlighted columns indicate precessing simulations. In all cases, it was possible to reconstruct at least three QNMs with a spatial mismatch of $\lesssim 10^{-5}$ with their associated spheroidal harmonics. In all cases, except the precessing and asymmetric 0013 simulation, the best-mapped mode is the fundamental $(2,2,0,+)$ QNM. The identity of the next best-mapped QNMs changes depending on the parameters of the progenitor binary. In all cases, the best three mapped QNMs are all prograde $n=0$ modes.
@@
The analysis in this section demonstrates that spatial mapping can effectively reconstruct the spatial structure of QNMs without explicitly providing mode-mixing coefficients. At first order, where PT makes clear predictions about the spatial structure, the technique offers another method for detecting QNMs and may, in analogy with the single- and multi-mode fits in Fig. 2, provide another way to determine the start time of the ringdown.
Figure 7: Leftmost panel: The real part of the normalized spatial reconstruction of the dominant QQNM with $\tilde{\alpha}=\tilde{\alpha}^{\prime}=(2,2,0,+)$ for the 0001 NR simulation. Other panels: The various predictions for the shape of this QQNM discussed in the main text. Prediction A clearly has qualitatively the wrong shape and is, therefore, discarded. The remaining B, C, and D predictions all have qualitatively similar shapes (from B(i) to D, the spatial mismatches with the reconstruction are $\mathcal{M}_{\tilde{\alpha}}=7.4\times 10^{-4}$, $\mathcal{M}_{\tilde{\alpha}}=2.9\times 10^{-4}$, $\mathcal{M}_{\tilde{\alpha}}=2.5\times 10^{-4}$, and $\mathcal{M}_{\tilde{\alpha}}=2.6\times 10^{-4}$) and are further investigated in Fig. 8.